OpenStudy (unklerhaukus):

\[\begin{align}a+b+c+d&=2\\ a^2+b^2+c^2+d^2&=30\\ a^3+b^3+c^3+d^3&=44\\ a^4+b^4+c^4+d^4&=354\end{align}\]

OpenStudy (unklerhaukus):

minor prize goes to the first to solve the system, major prize goes to the best method

OpenStudy (unklerhaukus):

clue; none of \(a,b,c,d\) are equal

OpenStudy (anonymous):

hmm nice clue LOL

OpenStudy (anonymous):

can we do either using substitution or elimination

OpenStudy (rajee_sam):

it is -1, 2, -3 and 4

OpenStudy (rajee_sam):

second equation is all squared and added. so 1 + 2 + 9 + 16 = 30 All numbers are different you said So put that idea in 1 and see 1 + 2 + 3 - 4 = 2 but when you apply the same to equation 3 the negativity becomes larger and adding all the cubes will turn it negative. so change the sign for 1 and 3 and 4 So the answer is -1,2,-3 and 4

OpenStudy (rajee_sam):

@UnkleRhaukus

OpenStudy (rajee_sam):

sorry squared and added is 1 + 4 + 9 + 16 = 30

OpenStudy (unklerhaukus):

Hey, @rajee_sam that is exactly the method i used !

OpenStudy (rajee_sam):

So where is my medal and fanning????

OpenStudy (rajee_sam):

what is my major prize???

OpenStudy (unklerhaukus):

\[\text{minor prize goes to}:\qquad\boxed{\stackrel{\color{goldenrod} {\ast\sim\star\sim\ast\sim\ast\sim\star\sim\ast}}{\color{purple}{\mathcal{RAJEE\_SAM}}}}\]

OpenStudy (rajee_sam):

minor prize?? you said major prize

OpenStudy (rajee_sam):

Ok first to solve I got it

OpenStudy (unklerhaukus):

major prize goes to the best method, i think there is another method , and i'd like to see it

OpenStudy (rajee_sam):

OK If there are no other?? What is the deadline for entries?? You think yours is not the best?? LOL

OpenStudy (anonymous):

well @UnkleRhaukus is there any other method

OpenStudy (unklerhaukus):

you tell me

OpenStudy (anonymous):

This is a good method for these type of problems. Use the given system to solve for these equations:\[a+b+c+d=p_1\]\[ab+ac+ad+bc+bd+cd=p_2\]\[abc+bcd+acd+abd=p_3\]\[abcd=p_4\]Then treat a,b,c and d as the four solutions to a quartic polynomial:\[x^4-p_1x^3+p_2x^2-p_3x+p_4=0\]The solutions you get will be a, b, c, and d. Here is an example with a system of three equations: http://openstudy.com/users/joemath314159#/updates/4e1002d70b8b56e555985df7